# Exploring the hardness of the ionizing radiation with the infrared softness diagram

### v1

@@ Abstract
Aims. After exploring a version of the infrared softness diagram to characterize the hardness of the incident ionizing radiation in star-forming regions, we exploit the availability of high-excitation lines in the same spectral regime to explore its potential use for studying the narrow-line regions of active galactic nuclei.
Methods. We adapted the IR softness diagram to consider very high-excitation lines, such as [Ne v] or [O iv]. The measured emission-line ratios were included as inputs for the code HCm-Teff-IR, in order to provide a quantification for both the ionization and the . The latter is sensitive to the spectral shape of the incident continuum in active galactic nuclei (AGN). We applied this to a large AGN sample including different spectral types with available Spitzer/IRS, Herschel/PACS and/or SOFIA/FIFI-LS mid-IR spectroscopic observations.
Results. The combination of the ([Ne ii]+[Ne iii])/[Ne v] and [O iii]/[O iv] emission line ratios is a robust proxy for the shape of the ionizing continuum in AGN. Given the difficulties to observe the [O iii] lines, an alternative based on the [S iii]+[S iv] lines can be used. The inclusion of very high-excitation lines in the input ratios for HCm-Teff-IR, to derive both and for the studied sample, points to a bimodal distribution of galaxies. One of the peaks is characterized by relatively harder values of around -1.4 in combination with low values for log around -2.4, while the other peak shows a softer -1.7 and high values of log -1.5. This result is consistent with the existence of two very marked AGN populations, one with a softer ionizing continuum, possibly dominated by a radiatively efficient accretion disk in bright Seyfert nuclei. In contrast, we observe a harder radiation field in low-luminosity AGN, where the accretion disk is expected to recede.
@@
The study and characterization of the nature of ionizing sources in galaxies has always been a recurrent area, given the diverse possibilities that can lead to the production of an emission-line spectrum typical of regions dominated by the emission from ionized gas. These can include episodes of massive star formation, stellar populations dominated by the presence of Hot Low-Mass Evolved Stars (HOLMES), shocks, or active galactic nuclei (AGN).
@@
Emission lines can provide valuable information on the gas properties and the sources responsible for the excitation and ionization of the elements producing them. In this context, given the observational difficulties found in directly studying the spectral regimes where the possible different sources can be distinguished (mainly X-rays and ultraviolet, UV), other diagnostics based on the relative fluxes of the lines emitted by the gas can be used instead, as these include strong recombination lines (RLs) and collisionally excited lines (CELs), more accessible as they are observed in the optical and the infrared (IR) ranges.
@@
The gas velocity dispersion derived from line widths provides important evidence of the mechanical energy supplied by different sources, and thus can be used to derive their nature. However, when sufficient spectral resolution is not available, alternative methods based on the line fluxes are required. In this case, there is a proliferation of diagnostic diagrams based on the flux of certain CELs relative to the strongest RLs, such as H or H. These include the well-known BPT diagrams (Baldwin et al., 1981), both empirically (e.g. Kauffmann et al., 2003) and theoretically (e.g. Kewley et al., 2006) calibrated, which are widely used to elucidate the nature of the ionizing source. On the other hand, as these diagnostic diagrams are based only on certain optical line fluxes, they are insufficient to produce a precise analysis of the incident spectral energy distribution (SED), and are most often used to establish a first-order selection in large samples prior to another kind of analysis, such as the determination of chemical abundances from these same emission lines. A more sophisticated approach to describe the nature and shape of the ionizing continuum in AGN was based on the analysis of several emission line ratios, both in the optical (Baldwin, 1975; Shields, 1978; Malkan & Sargent, 1982) and the IR range (Alexander et al., 1999, 2000; Meléndez et al., 2011). Nevertheless, most of these efforts were based on deterministic analyses using a relatively complex set of models, leading to ambiguous results on the shape of the ionizing continuum.
@@
An alternative approach to characterize the hardness of the radiation field in star-forming regions is the softness diagram (Pérez-Montero & Vílchez, 2009; Kumari et al., 2021), based on the flux ratios of emision lines from ionic species of consecutive ionization stages. This leverages the dependence of the so-called softness parameter (), defined by Vílchez & Pagel (1988), on the hardness of the ionizing radiation. In a previous paper of this series (Pérez-Montero et al. 2024, hereinafter Paper I), we analyzed a version of this diagram exclusively based on mid-IR lines, taking advantage of the large amount of data on emission-line galaxies available in this regime, such as the IDEOS catalog (Hernán-Caballero et al., 2016; Spoon et al., 2022), based on Spitzer/IRS, but also considering the upcoming new data from JWST.
@@
Among the various advantages that IR lines present compared to optical ones, their significantly lower dependence on extinction makes them especially appropriate for penetrating the large optical depths in galaxy dusty regions, particularly in deeply obscured objects such as Ultra-Luminous Infrared Galaxies (ULIRGs). Another major advantage of IR transitions is that their emissivities depend only marginally on the electron temperature, because the associated energy levels are much closer to the ground state. Additionally, this spectral regime presents transitions from highly ionized species that trace a more energetic spectral regime in the incident SED. This can reduce the uncertainty in the derivations affecting the optical tracers due to the contribution from low-excitation emission lines, such as [S ii] (Pérez-Montero et al., 2023).
@@
Several authors have proposed alternative parameters based on IR emission-lines from Ar, Ne, or S (Martín-Hernández et al., 2002; Morisset, 2004; Pérez-Montero & Vílchez, 2009), including the following expression examined in Paper I:
@@
We also described in Paper I the adaptation of the code HCm-Teff (Pérez-Montero et al., 2019b) to determine the equivalent effective temperature () and the ionization parameter () using IR transitions. This is based on a Bayesian-like comparison between the predictions of a comprehensive photoionization model grid and the appropriate IR emission-line ratios observed, as included in the definition of the softness diagram based on the parameter. This step is necessary to break the simultaneous dependence of on , , and metallicity (). When applied to the IDEOS sample, our results confirmed that no additional sources, harder than typical massive stars, were required to reproduce the emission-line ratios observed in all star-forming galaxies, including ULIRGs.
@@
In this work, we propose the study of softness parameters and diagrams based on IR lines to study the shape of the incident radiation field in AGN. Given the much harder SED expected in these objects, it is known that the more widely low-excitation lines in the optical are ineffective for this aim, so we explored the same parameter in the IR, assessing the use of other high-excitation emission lines in this range, such as [Ne v]14.3, 24.3 m or [O iv]25.9 m, whose presence has been predicted (Spinoglio & Malkan, 1992) and extensively confirmed in the literature for AGN (e.g. Genzel et al., 1998; Sturm et al., 2002; Meléndez et al., 2008; Armus et al., 2023).
@@
As in the case of for star-forming regions, we consider the index as a parameter that probes the shape of the incident SED in AGN. This index represents the slope of the power-law continuum between 2 keV and 2500 Å, and can be used to parametrize the hardness of the SED in each model. A high value indicates relatively strong X-ray emission compared to the UV range, which can highlight differences in the physics of the environment near the supermassive black hole (e.g. Tananbaum et al., 1979; Vignali et al., 2003). Fig. 1 shows the effect of changing the value in different double-peaked power laws, compared with the energy of the ions whose emission lines are typically observed in the mid-IR range, as well as with other SEDs typical of massive stars.
@@
The paper is organized as follows. In Section 2, we describe the observational samples of compiled mid- and far-IR emission lines in the sample of AGN. In Section 3, we describe the grids of photoionization models used to derive emission-line fluxes in the IR and build the different versions of the softness diagram. The most suitable form of this diagram for studying AGN is discussed in Section 4, along with a description of the adaptation of the HCm-Teff code for the IR range, adopting these models and lines for the analysis of AGN. In Section 5, we present and discuss our results from the application of this code to the different samples of objects. Finally, in Section 6, we summarize our results and present our conclusions.
@@ 2 Data samples
In this work, we compiled narrow nebular emission-line fluxes spectroscopically measured in the mid-IR in AGN by combining catalogs from Spitzer, Herschel, Akari, and SOFIA. The data were mainly obtained from three sources: 1) the Infrared Database of Extragalactic Observables111http://ideos.astro.cornell.edu (IDEOS Hernán-Caballero et al., 2016; Spoon et al., 2022) from Spitzer archive; 2) the AGN sample from Fernández-Ontiveros et al. (2016) and Spinoglio et al. (2022) combining Spitzer, Herschel and SOFIA; 3) the (U)LIRG catalog from Imanishi et al. (2010) observed with Akari.
@@
While the last two samples present an spectral classification of their AGN-dominated nature using both the optical (through BPT diagnostics) and IR ranges (through the presence of [Ne v] and [O iv] lines), this is not the case for the former one. In order to obtain a sample of AGN-dominated galaxies in the IDEOS sample, we used the strength (in terms of equivalent width) of the polycyclic aromatic hydrocarbon (PAH) emission at 6.2 m, which is expected to discriminate between star-formation and AGN activity (Puget & Leger, 1989; Armus et al., 2007; Hernán-Caballero et al., 2020). The selected sample verifies EW(PAH m, typical of galaxies dominated by AGN activity (Armus et al., 2007). Additionally, we required the detection of higher ionic species. Specifically, the selected sources have both [Ne v], either at 14m or 24m, and [O iv] detections in the IDEOS catalog, which are indicative of AGN activity (Armus et al., 2007). The subsequent analysis requires in addition [Ne ii] at 12.8 m, [Ne iii] at 15.6 m, and either [O iii] at 52 m or 88 m, or [S iv] at 10.5 m along with [S iii] at either 18.7 m or 33.5 m. Only emission line fluxes with a minimum signal-to-noise ratio (S/N) of 3 were used. After this selection to ensure the AGN-dominated nature of the selected objects, our final sample consists of 484 objects, considering as different objects different measurements taken at a distinct wavelength regime, spatial coverage or spectral resolution of the same galaxy. In any case, the number of duplicated objects (13) is very low and it does not affect the statistical significance of our results.
@@
As shown in Paper I for star-forming objects, the chemical content of the gas may be a relevant input parameter for the determination of the hardness using emission lines. Thus, we derived the total oxygen abundance in a small subsample of the resulting compiled AGN in order to better constrain the models used to characterize them. The determination of the oxygen abundance was done, as described in Pérez-Díaz et al. (2024), using a version of the HCm-IR code (Fernández-Ontiveros et al., 2021) adapted for its use for AGN (Pérez-Díaz et al., 2022). As described in Pérez-Díaz et al. (2022), the code only provides a solution for the derivation of the oxygen abundance of the gas-phase in AGN if at least one hydrogen recombination line is supplied. For this reason, available measurements of the hydrogen recombination line Brackett Akari/IRC (2.5-5 m, Imanishi et al. 2010), were added for some objects in the IDEOS catalog.
@@ 3 Description of the photoionization models
Several different grids of photoionization models were computed in order to be compared with the compiled observations, and provide a complete model database for the code HCm-Teff to obtain an estimate of the hardness of the incident radiation field in AGN.
@@
To this aim, we calculated grids of models similar to those described in Pérez-Montero et al. (2019a) and Pérez-Díaz et al. (2022), used for the calculation of chemical abundances in the narrow-line region (NLR) of AGN by HCm. These models use a composite SED as incident source, one component representing the Big Blue Bump peaking at 1 Ryd, and the other a power law with spectral index = -1, representing the non-thermal X-ray radiation in the 13.6 eV to 100 keV range (Elvis et al., 1994; Korista et al., 1997; Meléndez et al., 2011). As free parameter in the grids we computed models varying the values of the spectral index in the -2.0 to -0.8 range with a 0.2 step. Typical measurements of vary depending on the sample and the methods used. For instance, Miller et al. (2011) conducted a comprehensive study of in a large sample of AGN, finding values typically ranging from -1.5 to -1.1, depending on luminosity and redshift.
@@
HCm-IR and HCm-Teff use models computed with Cloudy v17.01222No significant differences were found between these models and test models run with Cloudy v23.01. (Ferland et al., 2017). The models span a range in total oxygen abundance from 12+log(O/H) = 8.1 to 9.0 in bins of 0.3 units, and log from -4.0 to -0.5 in bins of 0.25 units. These ranges cover the values derived in the different AGN samples analyzed using the code HCm (e.g. Pérez-Montero et al. 2019a; Dors et al. 2020; Pérez-Díaz et al. 2021, 2022; Pérez-Montero et al. 2023; Pérez-Díaz et al. 2024).
@@
All other chemical abundances were scaled to the solar proportions as described by Asplund et al. (2009), with the exception of N, which follows the empirical relation derived by Pérez-Montero et al. (2014). All models were calculated assuming a plane-parallel geometry, assuming a filling factor of 0.1 and a constant density of 500 cm-3, typical for these objects and below the critical density of the emission-lines involved in this study. A more thorough discussion on the effects of density in this models can be found in Pérez-Díaz et al. (2022, 2024). Two possible scenarios about the presence of dust grains mixed with the gas were considered in all models: one with a standard Milky Way proportion, and another way without dust grains in order to check the impact of the presence of dust in the results. In those models with dust, we did not consider depletion factors in the chemical abundances, so we always refer to the values in the gas-phase. In addition, the stopping criterion of each model is changed to consider either 99.9%, 98% or 2% of free electrons to explore how considering a matter-bounded geometry alters the emission-line ratios between low- and high-excitation lines. All these possibilities yield a total of 1 872 models in the final grid.
4 Determining the hardness in AGN from IR lines
4.1 A softness parameter in the mid-IR valid for AGN

### v2

@@ Abstract
Aims. After exploring a version of the infrared (IR) softness diagram to characterize the hardness of the incident ionizing radiation in star-forming regions, we exploit the availability of high-excitation lines in the same spectral regime to explore its potential use for studying the narrow-line regions of active galactic nuclei (AGNs).
Methods. We adapted the IR softness diagram to consider very high-excitation lines, such as [Ne v] or [O iv]. The measured emission-line ratios were included as inputs for the code HCm-Teff-IR in order to provide a quantification for the ionization and the . The latter is sensitive to the spectral shape of the incident continuum in AGNs. We applied this code to a large AGN sample that included different spectral types with available Spitzer/IRS, Herschel/PACS, and/or SOFIA/FIFI-LS mid-IR spectroscopic observations.
Results. The combination of the ([Ne ii]+[Ne iii])/[Ne v] and [O iii]/[O iv] emission line ratios is a robust proxy for the shape of the ionizing continuum in AGNs. Given the difficulties to observe the [O iii] lines, an alternative based on the [S iii]+[S iv] lines can be used. The inclusion of very high-excitation lines in the input ratios for HCm-Teff-IR to derive both and for the studied sample points to a bimodal distribution of galaxies. One of the peaks is characterized by relatively harder values of around -1.4 in combination with low values for log around -2.4, while the other peak shows a softer -1.7 and high values of log -1.5. This result is consistent with the existence of two very distinct AGN populations, where one has a softer ionizing continuum, possibly dominated by a radiatively efficient accretion disk in bright Seyfert nuclei. In contrast, we observed a harder radiation field in low-luminosity AGNs, where the accretion disk is expected to recede.
@@
The study and characterization of the nature of ionizing sources in galaxies has always been a recurrent area of research given the diverse possibilities that can lead to the production of an emission-line spectrum typical of regions dominated by the emission from ionized gas. These can include episodes of massive star formation, stellar populations dominated by the presence of hot low-mass evolved stars (HOLMES), shocks, or active galactic nuclei (AGNs).
@@
Emission lines can provide valuable information on the gas properties and the sources responsible for the excitation and ionization of the elements producing them. In this context, given the observational difficulties found in directly studying the spectral regimes where the possible different sources can be distinguished (mainly X-rays and ultraviolet), other diagnostics based on the relative fluxes of the lines emitted by the gas can be used instead, as they include strong recombination lines (RLs) and collisionally excited lines (CELs), which are more accessible since they are observed in the optical and the infrared (IR) ranges.
@@
The gas velocity dispersion derived from line widths provides important evidence of the mechanical energy supplied by different sources and thus can be used to derive their nature. However, when sufficient spectral resolution is not available, alternative methods based on the line fluxes are required. In this case, there is a proliferation of diagnostic diagrams based on the flux of certain CELs relative to the strongest RLs, such as H or H. These include the well-known Baldwin, Phillips, and Terlevich diagrams (BPT; Baldwin et al., 1981), both empirically (e.g., Kauffmann et al., 2003) and theoretically (e.g., Kewley et al., 2006) calibrated, which are widely used to elucidate the nature of the ionizing source. On the other hand, as these diagnostic diagrams are based only on certain optical line fluxes, they are insufficient to produce a precise analysis of the incident spectral energy distribution (SED) and are most often used to establish a first-order selection in large samples prior to another kind of analysis, such as the determination of chemical abundances from these same emission lines. A more sophisticated approach to describing the nature and shape of the ionizing continuum in AGNs is based on the analysis of several emission line ratios, both in the optical (Baldwin, 1975; Shields, 1978; Malkan & Sargent, 1982) and the IR range (Alexander et al., 1999, 2000; Meléndez et al., 2011). Nevertheless, most of these efforts are based on deterministic analyses using a relatively complex set of models, leading to ambiguous results on the shape of the ionizing continuum.
@@
An alternative approach to characterize the hardness of the radiation field in star-forming regions is the softness diagram (Pérez-Montero & Vílchez, 2009; Kumari et al., 2021), which is based on the flux ratios of emission lines from ionic species of consecutive ionization stages. This method leverages the dependence of the so-called softness parameter (), defined by Vílchez & Pagel (1988), on the hardness of the ionizing radiation. In a previous paper of this series (Pérez-Montero et al. 2024, hereinafter Paper I), we analyzed a version of this diagram exclusively based on mid-IR lines, taking advantage of the large amount of data on emission-line galaxies available in this regime, such as the IDEOS catalog (Hernán-Caballero et al., 2016; Spoon et al., 2022) based on Spitzer/IRS, but also considering the upcoming new data from JWST.
@@
Among the various advantages that IR lines present compared to optical ones is their significantly lower dependence on extinction, which makes them especially appropriate for penetrating the large optical depths in galaxy dusty regions, particularly in deeply obscured objects such as ultra-luminous infrared galaxies (ULIRGs). Another major advantage of IR transitions is that their emissivities depend only marginally on the electron temperature because the associated energy levels are much closer to the ground state. Additionally, this spectral regime presents transitions from highly ionized species that trace a more energetic spectral regime in the incident SED. This can reduce the uncertainty in the derivations affecting the optical tracers due to the contribution from low-excitation emission lines, such as [S ii] (Pérez-Montero et al., 2023).
@@
Several authors have proposed alternative parameters based on IR emission lines from Ar, Ne, or S (Martín-Hernández et al., 2002; Morisset, 2004; Pérez-Montero & Vílchez, 2009), including the following expression examined in Paper I:
@@
In Paper I, we also presented an adaptation of the HCm-Teff code (Pérez-Montero et al., 2019b), originally designed to determine the equivalent effective temperature () and the ionization parameter (), for use with IR transitions. This approach is based on a Bayesian-like comparison between the predictions of a comprehensive photoionization model grid and the appropriate IR emission-line ratios observed, as included in the definition of the softness diagram based on the parameter. This step is necessary to break the simultaneous dependence of on , , and metallicity (). When applied to the IDEOS sample, our results confirmed that no additional sources harder than typical massive stars were required to reproduce the emission-line ratios observed in all star-forming galaxies, including ULIRGs.
@@
In this work, we propose the study of softness parameters and diagrams based on IR lines in order to examine the shape of the incident radiation field in AGNs. Given the much harder SED expected in these objects, it is known that the more widely low-excitation lines in the optical are ineffective for this aim, so we explored the same parameter in the IR, assessing the use of other high-excitation emission lines in this range, such as [Ne v]14.3, 24.3 m or [O iv]25.9 m, whose presence has been predicted (Spinoglio & Malkan, 1992) and extensively confirmed in the literature for AGNs (e.g., Genzel et al., 1998; Sturm et al., 2002; Meléndez et al., 2008; Armus et al., 2023).
@@
As in the case of for star-forming regions, we consider the index as a parameter that probes the shape of the incident SED in AGNs. This index represents the slope of the power-law continuum between 2 keV and 2500 Å, and it can be used to parametrize the hardness of the SED in each model. A high value indicates relatively strong X-ray emission compared to the UV range, which can highlight differences in the physics of the environment near the supermassive black hole (e.g., Tananbaum et al., 1979; Vignali et al., 2003). Figure 1 shows the effect of changing the value in different double-peaked power laws compared with the energy of the ions whose emission lines are typically observed in the mid-IR range as well as with other SEDs typical of massive stars.
@@
The paper is organized as follows. In Sect. 2, we describe the observational samples of compiled mid- and far-IR emission lines in the sample of AGNs. In Sect. 3, we describe the grids of photoionization models used to derive emission-line fluxes in the IR and build the different versions of the softness diagram. The most suitable form of this diagram for studying AGNs is discussed in Sect. 4, which is accompanied by a description of the adaptation of the HCm-Teff code for the IR range. In Sect. 5, we present and discuss our results from the application of this code to the different samples of objects. Finally, in Sect. 6, we summarize our results and present our conclusions.
@@ 2 Data samples
In this work, we compiled narrow nebular emission-line fluxes spectroscopically measured in the mid-IR in AGNs by combining catalogs from Spitzer, Herschel, Akari, and SOFIA. The data were mainly obtained from three sources: 1) the Infrared Database of Extragalactic Observables111http://ideos.astro.cornell.edu (IDEOS Hernán-Caballero et al., 2016; Spoon et al., 2022) from Spitzer archive; 2) the AGN sample from Fernández-Ontiveros et al. (2016) and Spinoglio et al. (2022) combining Spitzer, Herschel and SOFIA; 3) the (U)LIRG catalog from Imanishi et al. (2010) observed with Akari.
@@
While the last two samples present an spectral classification of their AGN-dominated nature using both the optical (through BPT diagnostics) and IR ranges (through the presence of [Ne v] and [O iv] lines), this is not the case for the former one. In order to obtain a sample of AGN-dominated galaxies in the IDEOS sample, we used the strength (in terms of equivalent width; EW) of the polycyclic aromatic hydrocarbon (PAH) emission at 6.2 m, which is expected to discriminate between star-formation and AGN activity (Puget & Leger, 1989; Armus et al., 2007; Hernán-Caballero et al., 2020). The selected sample verifies a EW(PAH m, which is typical of galaxies dominated by AGN activity (Armus et al., 2007). Additionally, we required the detection of higher ionic species. Specifically, the selected sources have both [Ne v], either at 14m or 24m, and [O iv] detections in the IDEOS catalog, which are indicative of AGN activity (Armus et al., 2007). The subsequent analysis requires in addition [Ne ii] at 12.8 m, [Ne iii] at 15.6 m, and either [O iii] at 52 m or 88 m, or [S iv] at 10.5 m along with [S iii] at either 18.7 m or 33.5 m. Only emission line fluxes with a minimum signal-to-noise ratio (S/N) of 3 were used. After this selection to ensure the AGN-dominated nature of the selected objects, our final sample consists of 484 objects, considering as different objects different measurements taken at a distinct wavelength regime, spatial coverage or spectral resolution of the same galaxy. In any case, the number of duplicated objects (13) is very low and it does not affect the statistical significance of our results.
@@
As shown in Paper I for star-forming objects, the chemical content of the gas may be a relevant input parameter for the determination of the hardness using emission lines. Thus, we derived the total oxygen abundance in a small subsample of the resulting compiled AGNs in order to better constrain the models used to characterize them. The determination of the oxygen abundance was done, as described in Pérez-Díaz et al. (2024), using a version of the HCm-IR code (Fernández-Ontiveros et al., 2021) adapted for its use for AGNs (Pérez-Díaz et al., 2022). As described in Pérez-Díaz et al. (2022), the code only provides a solution for the derivation of the oxygen abundance of the gas-phase in AGNs if at least one hydrogen recombination line is supplied. For this reason, available measurements of the hydrogen recombination line Brackett Akari/IRC (2.5-5 m, Imanishi et al. 2010), were added for some objects in the IDEOS catalog.
@@ 3 Description of the photoionization models
Several different grids of photoionization models were computed in order to be compared with the compiled observations, and provide a complete model database for the code HCm-Teff to obtain an estimate of the hardness of the incident radiation field in AGNs.
@@
To this aim, we calculated grids of models similar to those described in Pérez-Montero et al. (2019a) and Pérez-Díaz et al. (2022), used for the calculation of chemical abundances in the narrow-line region (NLR) of AGNs by HCm. These models use a composite SED as incident source, one component representing the Big Blue Bump peaking at 1 Ryd, and the other a power law with spectral index = -1, representing the non-thermal X-ray radiation in the 13.6 eV to 100 keV range (Elvis et al., 1994; Korista et al., 1997; Meléndez et al., 2011). As free parameter in the grids we computed models varying the values of the spectral index in the -2.0 to -0.8 range with a 0.2 step. Typical measurements of vary depending on the sample and the methods used. For instance, Miller et al. (2011) conducted a comprehensive study of in a large sample of AGNs, finding values typically ranging from -1.5 to -1.1, depending on luminosity and redshift.
@@
HCm-IR and HCm-Teff use models computed with Cloudy v17.01222No significant differences were found between these models and test models run with Cloudy v23.01. (Ferland et al., 2017). The models span a range in total oxygen abundance from 12+log(O/H) = 8.1 to 9.0 in bins of 0.3 units, and log from -4.0 to -0.5 in bins of 0.25 units. These ranges cover the values derived in the different AGN samples analyzed using the code HCm (e.g., Pérez-Montero et al. 2019a; Dors et al. 2020; Pérez-Díaz et al. 2021, 2022; Pérez-Montero et al. 2023; Pérez-Díaz et al. 2024).
@@
All other chemical abundances were scaled to the solar proportions as described by Asplund et al. (2009), with the exception of N, which follows the empirical relation derived by Pérez-Montero et al. (2014). All models were calculated assuming a plane-parallel geometry, assuming a filling factor of 0.1 and a constant density of 500 cm-3, typical for these objects and below the critical density of the emission lines involved in this study. A more thorough discussion on the effects of density in this models can be found in Pérez-Díaz et al. (2022, 2024). Two possible scenarios about the presence of dust grains mixed with the gas were considered in all models: one with a standard Milky Way proportion, and another way without dust grains in order to check the impact of the presence of dust in the results. In those models with dust, we did not consider depletion factors in the chemical abundances, so we always refer to the values in the gas-phase. In addition, the stopping criterion of each model is changed to consider either 99.9%, 98% or 2% of free electrons to explore how considering a matter-bounded geometry alters the emission-line ratios between low- and high-excitation lines. All these possibilities yield a total of 1 872 models in the final grid.
4 Determining the hardness in AGNs from IR lines
4.1 A softness parameter in the mid-IR valid for AGNs